A strain guage has a resistance of 120.2 ohms at 20oC. Calculate its resistance at 30oC.


1
Expert's answer
2021-08-16T16:35:07-0400

"R_0=120 ohm \\T_1=20°c\\T_2=30°c"

"R_1=?"

"R_1=R_0(1+\alpha(T_2-T_1))"

"\alpha=8\times10^{-7}\Omega/\Omega°c"

Put value


"R_1=120\times(1+8\times10^{-7}(30-20))"

"R_1=120\times(1+8\times10^{-7}\times(10))"

"R_1=120.61\Omega"